OpenStudy (joftheworld):

@precal can you help me plot three points from a few linear equations ? its difficult because I need to find whole numbers for the points... the first linear equation I have to plot is -3y=5x-7

OpenStudy (precal):

just divide -3 to both sides and what do you get?

OpenStudy (joftheworld):

do I divide 5x by -3y? or -7 ?

OpenStudy (precal):

yes the whole right side and left side, what do you get?

OpenStudy (joftheworld):

I don't get it ..

OpenStudy (precal):

the whole point is to get this equation into what we call point-slope form or just isolate y by itself. This will help you quickly graph the equation and find points from the line.

OpenStudy (joftheworld):

can you help me get the first point ?

OpenStudy (precal):

let's rewrite the equation first, there is much to learn by doing this: I have a purpose to what I am teaching you. Remember you asked for my help. I am here to help.

OpenStudy (joftheworld):

im so confused ...

OpenStudy (precal):

ok let's take another approach, if you just want three points just sub any value into your equation

OpenStudy (precal):

-3y=5x-7 let's just pick any points shall we

OpenStudy (precal):

if x=0 what is y?

OpenStudy (joftheworld):

I tried plugging in 0 for x.. did I mention I need whole numbers ? I cant use fractions or decimals in the points.. because its an online course..

OpenStudy (precal):

-3y=5x-7 -3y=5(0)-7 -3y=-7 y=7/3 ok here is your first point (0, 7/3)

OpenStudy (precal):

really fractions and decimals should not matter, even on an online course.....this is math we are not restricted to whole numbers only

OpenStudy (precal):

if you just want whole numbers, use a graphing calculator and look up the values in the table.

OpenStudy (joftheworld):

it matters in this online course.. sadly the graph wont let me plot in between numbers..

OpenStudy (precal):

no there are actually whole numbers on this line, don't you have a graphing calculator?

OpenStudy (joftheworld):

no I do not..